# Investigating ReLoRA: Effects on the Learning Dynamics of Small Language Models

### v1

@@ Abstract
Parameter-efficient methods such as LoRA have revolutionised the fine-tuning of LLMs. Still, their extension to pretraining via ReLoRA is less well understood, especially for small language models (SLMs), which offer lower computational and environmental costs. This work is the first systematic study of ReLoRA in SLMs (11M–66M parameters), evaluating both performance and learning dynamics. Through ablation experiments, we find that ReLoRA generally performs worse than standard training on loss, Paloma perplexity and BLiMP, with the gap widening for the larger models. Further analysis of the learning dynamics of the models indicates that ReLoRA reinforces the rank deficiencies found in smaller models. These results indicate that low-rank update strategies may not transfer easily to SLM pretraining, highlighting the need for more research in the low-compute regime.
@@ 1 Introduction
Contemporary work on language modelling, particularly in industry, has continually prioritised ever-larger scales (Chowdhery et al., 2023; Grattafiori et al., 2024; OpenAI, 2024). Nevertheless, these impressive performance gains come at a significant cost. For instance, recent findings indicate that training a 13-billion-parameter model with a Llama-style architecture (Touvron et al., 2023) uses an equivalent amount of energy to what an average US home uses in nine years and seven months (Morrison et al., 2024). Furthermore, language models (LMs) can perpetuate pernicious social biases through representational and allocational harms (Bender et al., 2021; Gallegos et al., 2024), and larger models can leak personal information through their memorisation patterns (Huang et al., 2022). Therefore, spurred by the push towards Green AI (Schwartz et al., 2020), which centres efficiency as an essential evaluation metric, research into small language models (SLMs) is vital. However, the performance of LMs is extraordinarily dependent on scale (Kaplan et al., 2020), and SLMs consistently underperform their larger counterparts.
@@
Regardless of size, LMs are challenging to pretrain (Mielke et al., 2019; Ettinger, 2020), and fine-tuning can be similarly resource-intensive for large models (Touvron et al., 2023). Parameter-efficient fine-tuning (PEFT) addresses this by updating only critical layers, thus reducing computational overheads (Houlsby et al., 2019; Ding et al., 2023; Han et al., 2024). One such method, LoRA (Hu et al., 2022), has been widely used due to its effectiveness and applicability to any model with matrix computations, enabling straightforward fine-tuning of even the largest models (Blattmann et al., 2023; Dettmers et al., 2023; Fomenko et al., 2024). LoRA introduces low-rank adaptation, where the pretrained model weights are frozen, and fine-tuning is performed with trainable rank decomposition matrices at each layer of the architecture, thus significantly reducing the number of trainable parameters.
@@
ReLoRA, proposed by Lialin et al. (2023), aims to leverage LoRA’s overwhelming success in low-rank adaptation for fine-tuning to advance efficient pretraining. It does this by injecting low-rank LoRA-style matrices into an LM, then repeatedly merging and reinitialising them throughout the training process, thereby parametrising gradient updates through the merge operations at each restart. The method is tested on models with 60M to 1.3B parameters, improving training speeds and reducing GPU memory footprints, especially at the higher end. However, research for models even smaller than this is limited. Small transformer models suffer from rank deficiencies in their weight matrices, limiting the subspace they can explore during training (Diehl Martinez et al., 2024). This makes them a sensitive setting in which to detect whether ReLoRA’s periodic low-rank merges either expand the accessible parameter subspace or instead erase critical structures. Their reduced scale amplifies both potential gains and drawbacks, providing a fast, low-cost platform for probing the method’s dynamics before scaling up.
Given this, we explore whether ReLoRA is effective in the small/tiny language modelling domain. In other words, does ReLoRA boost performance or drag it down?
[…]
@@
thumbsup Boost: By strategically resetting and merging LoRA matrices, ReLoRA aggregates multiple low-rank steps, which may widen those bottlenecks and enable smaller models to capture more complex patterns
[…]
@@
thumbsdown Drag: However, ReLoRA may also have the opposite effect: further reducing the (already limited) representational capacity of SLMs.
@@
[…]
    thumbsdown Drag: Evidence of performance degradation and training instability. Across loss, Paloma perplexity, and BLiMP evaluation, ReLoRA underperforms its full-rank baseline, with widening gaps as model size increases from the tiny to small scales. Our findings suggest that low-rank strategies do not trivially transfer to SLMs and thus motivate hybrid or adaptive-rank approaches for future low-resource model training.
@@ 3 Background: Rank Dynamics
Relatedly, studies of rank dynamics in transformers have shown two interesting effects; first, weight matrices tend to increase in effective rank gradually over training (Boix-Adsera et al., 2023), expanding their representational capacity, second, at the same tine “rank collapse” at initialisation can severely limit early gradient flow in key and query projections (Noci et al., 2022). These findings suggest that the ability to build rank over time is important for model quality. How ReLoRA intersects with these findings in the context of small language models is an open question and the study of this paper. We hypothesise two possible effects.
@@ 5.2.1 Preparation
Attention heads can be conceptualised as being composed of two essentially independent calculations: a ‘Query-Key’ circuit, which calculates the attention pattern, and an ‘Output-Value’ (OV) circuit, which computes the amount a given token affects output logits (if it is attended to) (Elhage et al., 2021). The OV circuit is calculated by taking the matrix product of the Output and Value matrices. OV circuits, despite being parameterised separately, may be considered as individual, low-rank matrices.
Learning dynamics.
Following this, learning dynamics analysis is performed on activations and parameters of the attention and feed-forward layers that ‘write’ to the residual stream, namely the OV circuit and SwiGlu $W_{2}$ layers.
@@ 5.2.2 Proportional effective rank
Given what we know about the rank deficiencies in smaller models, we wish to probe how they behave and develop. To do this, we use PER (Diehl Martinez et al., 2024), a metric which enables size-agnostic comparison of the effective rank of parameter weight matrices. We measure PER on the weight matrices in the feed-forward and attention layers that parameterise the ‘writes’ back to the residual stream. PER is given by Equation 4 below.
@@ 5.2.4 Findings from PER and CN
[…]
@@
[…]
@@
This behaviour is compounded by the anisotropy of SLMs, which refers to the phenomenon where the internal representations of tokens are unevenly distributed and highly clustered within the representational space (Godey et al., 2024). As a result, even small changes in input can cause a token embedding to shift into a sparse or differently clustered region, thereby amplifying differences in the model’s output.
@@ 6.1 Findings
[…]
    thumbs down Drag: ReLoRA degrades pretraining performance in SLMs. Across loss, Paloma perplexity, and BLiMP evaluation, ReLoRA consistently underperforms conventional full-rank training. The performance gap is minor for tiny models but grows substantially as model size increases to the small scale.
@@
[…]
    thumbs down Drag: ReLoRA exacerbates low-rank bottlenecks and training instability. Learning dynamics analysis indicates that ReLoRA leads to reduced PERs of the models’ parameters and gradient updates. Alongside this, ReLoRA induces highly ill-conditioned gradient updates that increase the models’ susceptibility to numerical errors. Both of these exacerbate existing issues caused by the inherent anisotropy of SLM representations.
@@ 6.2 Implications
This paper highlights a principal limitation of low-rank pretraining for SLMs, suggesting that for a parameter-efficient training method to succeed in the low-compute domain, it likely must preserve higher-rank update spaces or adapt to models’ intrinsic dimensionalities.
Future investigations should focus on hybrid approaches. For instance, coupling low-rank adapters with selective full-rank updates or utilising dynamic rank adaptation techniques (such as DyLoRA (Valipour et al., 2023)) to minimise representational losses. Our results aim to inform the design of future low-resource LM training algorithms, emphasising that efficiency gains cannot come at the cost of substantial expressivity and performance loss.

### v2

@@ Abstract
Parameter-efficient methods like LoRA have revolutionised large language model (LLM) fine-tuning. ReLoRA extends this idea to pretraining by repeatedly merging and reinitialising low-rank adapters, increasing cumulative rank while keeping updates cheap. This aligns well with observations that high-capacity models learn through locally low-rank trajectories that expand over time. By contrast, recent work suggests that small language models (SLMs) exhibit rank deficiencies and under-utilise their available dimensionality. This raises a natural question: can ReLoRA’s rank-expanding update rule steer SLMs toward healthier learning dynamics, mitigating rank bottlenecks in a capacity-constrained regime? We argue SLMs are an ideal testbed: they train quickly, enable controlled ablations, and make rank phenomena more measurable. We present the first systematic study of ReLoRA in SLMs (11M–66M parameters), evaluating both performance and learning dynamics. Across loss, Paloma perplexity, and BLiMP, we find that ReLoRA underperforms full-rank training, with gaps widening at larger scales. Analysis of proportional effective rank and condition numbers shows that ReLoRA amplifies existing rank deficiencies and induces ill-conditioned updates early in training. Our results suggest that while ReLoRA’s merge-and-restart strategy can expand ranks in larger models, it does not straightforwardly translate to capacity-limited SLMs, motivating adaptive-rank or hybrid-rank approaches for low-compute pretraining.
@@ 1 Introduction
Contemporary work on language modelling has continually prioritised ever-larger scales, delivering remarkable capability gains (Chowdhery et al., 2023; Grattafiori et al., 2024; OpenAI, 2024). However, the computational and environmental costs of large-scale language modelling are substantial (Chowdhery et al., 2023; Grattafiori et al., 2024; Morrison et al., 2024; OpenAI, 2024). Small language models (SLMs) offer a complementary path: they are cheap to train and deploy, easier to study, and attractive for settings where safety, privacy, or energy constraints dominate (Schwartz et al., 2020). However, SLMs lag in quality due to tight capacity limits and brittle optimisation (Mielke et al., 2019; Ettinger, 2020; Kaplan et al., 2020).
A rank-centric view of learning.
A growing body of work frames transformer learning through the lens of the rank of model weights. In high-capacity models, training often proceeds via low-rank updates that gradually expand effective rank (Aghajanyan et al., 2021; Boix-Adsera et al., 2023). Conversely, SLMs show rank deficiencies and anisotropic representations that restrict gradient flow and under-utilise model dimensionality (Noci et al., 2022; Diehl Martinez et al., 2024; Godey et al., 2024a). From this perspective, methods that shape the rank profile of updates could directly influence optimisation quality and final capability.
LoRA.
@@
LoRA (Hu et al., 2022) has been widely used due to its effectiveness and applicability to any model with matrix computations, enabling straightforward fine-tuning of even the largest models (Blattmann et al., 2023; Dettmers et al., 2023; Fomenko et al., 2024). LoRA introduces low-rank adaptation, where the pretrained model weights are frozen, and fine-tuning is performed with trainable rank decomposition matrices at each layer of the architecture, thus significantly reducing the number of trainable parameters.
ReLoRA.
@@
ReLoRA, proposed by Lialin et al. (2023), aims to leverage LoRA’s overwhelming success in low-rank adaptation for fine-tuning to advance efficient pretraining. It does this by injecting low-rank LoRA-style matrices into a language model, then repeatedly merging and reinitialising them throughout the training process, thereby parametrising gradient updates through the merge operations at each restart. The method is tested on models with 60M to 1.3B parameters, improving training speeds and reducing GPU memory footprints, especially at the higher end. However, research for models even smaller than this is limited.
Why ReLoRA might help SLMs.
Small transformer models suffer from rank deficiencies in their weight matrices, limiting the subspace they can explore during training (Diehl Martinez et al., 2024). ReLoRA (Lialin et al., 2023) periodically merges low-rank adapters into the base weights and reinitialises them. Summing distinct low-rank updates increases the rank of the cumulative update, potentially widening the subspace explored over training. If SLMs struggle because their updates remain trapped in low-rank bottlenecks, then an explicit rank-expansion mechanism could steer them toward healthier dynamics, improving sample efficiency and language competence in the low compute regime. This provides the intuition behind our methodology.
Why test this in SLMs?
SLMs are fast to train, enabling careful ablations across model structure alongside detailed diagnostics such as effective rank and condition number that would otherwise be prohibitively costly in billion-parameter models. Their capacity limits also amplify rank phenomena, making both benefits and failure modes easier to detect. Therefore, SLMs are a sensitive and economical sandbox for understanding how rank-structured updates shape learning.
This work: research question.
In this work, we ask whether ReLoRA’s rank-expanding merge-and-restart methodology actually helps in the capacity-constrained regime. In other words, does ReLoRA boost performance or drag it down? The case for either can be summarised as follows:
@@
[Uncaptioned image] Boost: By strategically resetting and merging LoRA matrices, ReLoRA aggregates multiple low-rank steps, which may widen those bottlenecks and enable smaller models to capture more complex patterns
@@
[Uncaptioned image] Drag: However, ReLoRA may also have the opposite effect: further reducing the (already limited) representational capacity of SLMs.
This work: outline of methodology and results.
We extend a Llama-style SLM (Touvron et al., 2023; Diehl Martinez, 2025) with ReLoRA and run matched ablations at 11M and 66M parameters. We evaluate training loss, Paloma perplexity (Magnusson et al., 2024), and BLiMP (Warstadt et al., 2020), and additionally probe learning dynamics via proportional effective rank (PER) and condition numbers of both weights and updates. Contrary to the motivating intuition, ReLoRA does not boost SLMs: it underperforms full-rank training and reinforces rank deficiencies, with early training marked by highly ill-conditioned updates. We thus conclude that the mechanism that helps larger models does not straightforwardly transfer to SLMs.
This work: contributions.
@@
    [Uncaptioned image] Drag: Evidence of performance degradation and training instability. Across loss, Paloma perplexity, and BLiMP evaluation, ReLoRA underperforms its full-rank baseline, with widening gaps as model size increases from the tiny to small scales. Our findings suggest that low-rank strategies do not trivially transfer to SLMs and thus motivate hybrid or adaptive-rank approaches for future low-resource model training.
@@ 3 Background: Rank Dynamics
Relatedly, studies of rank dynamics in transformers have shown two interesting effects: first, weight matrices tend to increase in effective rank gradually over training (Boix-Adsera et al., 2023), expanding their representational capacity; second, at the same time, “rank collapse” at initialisation can severely limit early gradient flow in key and query projections (Noci et al., 2022). These findings suggest that the ability to build rank over time is important for model quality. How ReLoRA intersects with these findings in the context of small language models is an open question and the subject of this paper. We hypothesise two possible effects.
@@ 5.2.1 Preparation
Attention heads can be conceptualised as being composed of two essentially independent calculations: a ‘Query-Key’ circuit, which calculates the attention pattern, and an ‘Output-Value’ (OV) circuit, which computes the amount a given token affects output logits (if it is attended to) (Elhage et al., 2021). The OV circuit is calculated by taking the matrix product of the Output and Value matrices. OV circuits, despite being parameterised separately, may be considered as individual, low-rank matrices. We focus on the OV circuit as it ‘writes’ to the residual stream, a part of the model that has been shown to suffer from performance bottlenecks via the output representations of the model (Godey et al., 2024b).
Learning dynamics & investigated parameters.
Following this, learning dynamics analysis is performed on certain activations and parameters of the models’ attention and feed-forward layers. Specifically, we choose to analyse activations and parameters that ‘write’ to the residual stream, namely the OV circuit and SwiGLU $W_{2}$ layers (Elhage et al., 2021). Therefore, their behaviour has a salient influence on the evolution of internal representations across layers. This makes them particularly valuable probes for rank bottlenecks and deficiencies.
@@ 5.2.2 Proportional effective rank
Given what we know about the rank deficiencies in smaller models, we aim to investigate how they behave and evolve. To achieve this, we utilise PER (Diehl Martinez et al., 2024), a metric that enables size-agnostic comparison of the effective rank of parameter weight matrices. We measure PER on the weight matrices in the feed-forward and attention layers that parameterise the ‘writes’ back to the residual stream. PER is given by Equation 4 below.
@@ 5.2.4 Findings from PER and CN
[Uncaptioned image] Low PER values drag ReLoRA’s performance down.
@@
[Uncaptioned image] pico-relora is considerably more sensitive to fluctuations in inputs.
@@
This behaviour is compounded by the anisotropy of SLMs, which refers to the phenomenon where the internal representations of tokens are unevenly distributed and highly clustered within the representational space (Godey et al., 2024a). As a result, even small changes in input can cause a token embedding to shift into a sparse or differently clustered region, thereby amplifying differences in the model’s output.
@@ 6.1 Findings
    [Uncaptioned image] Drag: ReLoRA degrades pretraining performance in SLMs. Across loss, Paloma perplexity, and BLiMP evaluation, ReLoRA consistently underperforms conventional full-rank training. The performance gap is minor for tiny models but grows substantially as model size increases to the small scale. However, this might be caused by minimal parameter differences between the two tiny models, leading to a surprising performance loss (compared to the findings of Lialin et al. (2023) on large models) as model scale increases. One possible explanation is that the tiny model’s parameterisation leaves little difference between ReLoRA and full-rank updates, therefore leading to a lower-than-expected drop in performance.
@@
    [Uncaptioned image] Drag: ReLoRA exacerbates low-rank bottlenecks and training instability. Learning dynamics analysis indicates that ReLoRA leads to reduced PERs of the models’ parameters and gradient updates. Alongside this, ReLoRA induces highly ill-conditioned gradient updates that increase the models’ susceptibility to numerical errors. Both of these exacerbate existing issues caused by the inherent anisotropy of SLM representations.
@@ 6.2 Implications
This paper highlights a principal limitation of low-rank pretraining for SLMs, suggesting that for a parameter-efficient training method to succeed in the low-compute domain, it likely must preserve higher-rank update spaces or adapt to models’ intrinsic dimensionalities. Furthermore, our results prompt reflection on whether parameter-efficient training methods are necessary for SLMs at all. Unlike SOTA multi-billion-parameter models, SLMs can be trained on a single modern GPU (and often even on consumer hardware), raising the question of whether parameter-efficiency provides any benefits in this regime.
While these are negative results, they are informative in highlighting the boundaries at which parameter-efficient pretraining breaks down. Our study thus highlights where new methods are most needed.
Future investigations should therefore focus on novel hybrid approaches. For instance, coupling low-rank adapters with selective full-rank updates or utilising dynamic rank adaptation techniques (such as DyLoRA (Valipour et al., 2023)) to minimise representational losses. Our results aim to inform the design of future low-resource LM training algorithms, emphasising that efficiency gains cannot come at the cost of substantial expressivity and performance loss.